1. While keeping equilibrium reactions and solubility reactions in mind, what connection would you assume to exist between an acidity constant and how acidic a solution would be?
1
Expert's answer
2020-01-17T07:01:37-0500

Exist between an acidity constant and how acidic a solution is Henderson–Hasselbalch equation: "{\displaystyle \mathrm {pH} =\mathrm {p} K_{\mathrm {a} }+\log \mathrm {\frac {[A^{-}]}{[HA]}} }"